OpenStudy (anonymous):

Adam walked up to Bob and gave him 1/4 of his apples. He then came up to Charlotte and gave her 1/2 of the apples he had left. Charlotte then went to Bob to give him 1/3 of the apples she had. Bob then decided to go back to Adam to give him 2 of the apples he had. How many apples does each person have at the end and how many apples did Adam lose?

OpenStudy (chihiroasleaf):

Let A be Adam's apple, B be Bob's apple, C be Charlotte's apple from the first sentence, B_1 = B + (1/4) A from the second sentence C_1 = C + 1/2 ((3/4) A) then, B = B_1 + (1/3) C_1 A_1 = (1 - (1/4) - (3/8)) + 2 I think, you should solve this equations, if I don't do any mistake in make the model

OpenStudy (anonymous):

I don't know, I just made this up.

OpenStudy (anonymous):

But I'll give you a medal for trying :3